A gas has a volume of 19.56 L at a temperature of 183 K. What will its volume be at a temperature of 327 K?


Item 2 is unpinned. Click to pin.

A gas has is cooled from 106.43 oC to 43.15 oC. If its final volume is 4.79 L, what was its initial volume?


A gas is heated and its volume increases from 5.02 to 18.16 L. If the final temperature is 118 oC what was its initial temperature in Kelvin?


The volume of a gas increase from 3.43 to 8.80 L. If the initial temperature was 107 K, what is the final temperature?


If a gas has a volume of 22.30 L at 481 K, at what temperature will it have a volume of 7.83 L?


1
Expert's answer
2021-05-19T02:27:10-0400

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


"\frac{19.56}{183}=\frac{V_2}{327}"


"V_2=34.95l"





"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


"\frac{V_1}{379.43}=\frac{4.79}{316.15}"


"V_1=5.75l"






"\frac{5.02}{T_1}=\frac{18.16}{391}"


"T_1=108.08K"